OpenStudy (anonymous):

What is the value of n in the equation five-sixths equals n divided by five plus one-half? n equals one-fifteenth n equals three-fifths n equals one and two-thirds n equals 5

OpenStudy (yanasidlinskiy):

Again, 5/6 = n/5 + 1/2 -- subtract 1/2 from both sides 5/6 - 1/2 = n/5 5/6 - 3/6 = n/5 1/3 = n/5 -- multiply both sides by 5 1/3(5) = n 5/3 = n 1 2/3 = n

OpenStudy (anonymous):

thx

OpenStudy (yanasidlinskiy):

No problem!:)